3x + 2y = 2 -2x + y = 8
Answer by algebrahouse.com(1659)

3x + 2y = 2 -----> 3x + 2y = 2 {top equation stays the same}
-2x + y = 8 -----> 4x - 2y = -16 {multiplied bottom equation by -2}

7x = -14 {added the two new equations together}
x = -2 {divided both sides by 7}

3x + 2y = 2 {original top equation}
3(-2) + 2y = 2 {substituted -2, in for x, into top equation}
-6 + 2y = 2 {multiplied 3 by -2}
2y = 8 {added 6 to both sides}
y = 4 {divided both sides by 2}

x = -2 and y = 4

For more help from me, visit: www.algebrahouse.com